OpenStudy (anonymous):

2|x-1| < x^2 answers are: x < -1 - sqrt(3) or x > -1+sqrt(3) I'm having trouble coming up with these solutions. Could someone help me with the steps?

myininaya (myininaya):

Did you try looking at 2|x-1|=x^2 first When we have |f(x)|=a, we try to solve this by doing f(x)=-a of f(x)=a Now keep in mind that a needs to be positive or 0. Guess what? x^2 is a positive number or 0.

myininaya (myininaya):

I'm tell you to solve the following 2(x-1)=x^2 or 2(x-1)=-x^2

OpenStudy (anonymous):

I've got it now. I thought that I tried that before, but I think I was just making so many mistakes that I became too sloppy and frustrated

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

getting rid of the inequalities made it a lot simpler.

OpenStudy (anonymous):

how would you know how to replace the inequalities back into the answer?

OpenStudy (anonymous):

analytically, I mean

myininaya (myininaya):

One of the equations you solve, you will get a complex answer. The other equation you solve, will give you two real solutions. You should see x=-1+sqrt(3) or x=-1-sqrt(3). You can test intervals to see where we have 2|x-1|<x^2 

myininaya (myininaya):

 2|-10-1|<10^2 True/False 2|0-1|<0^2 True/False 2|10-1|<10^2 True/False